# Scalable Population Training for Zero-Shot Coordination

### v1

@@ Abstract
Zero-shot coordination(ZSC) has become a hot topic in reinforcement learning research recently. It focuses on the generalization ability of agents, requiring them to coordinate well with collaborators that are not seen before without any fine-tuning. Population-based training has been proven to provide good zero-shot coordination performance; nevertheless, existing methods are limited by computational resources, mainly focusing on optimizing diversity in small populations while neglecting the potential performance gains from scaling population size. To address this issue, this paper proposes the Scalable Population Training (ScaPT), an efficient training framework comprising two key components: a meta-agent that efficiently realizes a population by selectively sharing parameters across agents, and a mutual information regularizer that guarantees population diversity. To empirically validate the effectiveness of ScaPT, this paper evaluates it along with representational frameworks in Hanabi and confirms its superiority.
@@ Introduction
As MARL continues to achieve remarkable success in a wide range of domains, its scalability is increasingly hindered by the greedy growth of computational resource demands, in the scenarios like agent MOBA games (gao2023towards), robotic soccer (zhu2024dynamic), and intelligent manufacturing systems (do2025heterogeneous) owing to the demand for training a diverse set of agents. These challenges highlight the need for more scalable and resource-efficient MARL frameworks.
@@
We address this problem in the context of Zero-Shot Coordination (ZSC)—a setting where agents must collaborate with unseen partners. While self-play (lowe2017multi) is widely used for training cooperation, it often leads to overfitting, limiting agents’ ability to generalize. To overcome this, hu2020other introduced the ZSC problem, where agents are evaluated on their ability to coordinate without prior exposure to their partners. Population-based training (jaderberg2017population) offers a promising solution by exposing agents to a diverse set of behaviors during training, thus encouraging more generalizable coordination strategies (charakorn2022generating). In principle, if the training population covers all potential partner strategies, ZSC would be effectively solved. However, approaching this ideal requires training a large and diverse population of agents, which is itself constrained by limited computational resources, hindering the development and scalability of such methods.
Figure 1: The diagram of different training paradigms.
@@
Recent methods focus on improving Zero-Shot Coordination (ZSC) performance by enhancing population diversity. For example, Trajectory Diversity for Zero-Shot Coordination (lupu2021trajectory) maximizes trajectory-level diversity to foster generalizable strategies. Similarly, Maximum Entropy Population-Based Training (zhao2023maximum) uses entropy maximization to broaden behavioral coverage and improve human-AI coordination. Heterogeneous Multi-Agent Zero-Shot Coordination by Coevolution (xue2024heterogeneous) tackles real-world heterogeneity by co-evolving two distinct populations through iterative pairing and selection, enabling robust cooperation with diverse partners.
@@
However, constrained by limited computational resources, existing studies typically maintain relatively small population sizes and attempt to improve the zero-shot coordination performance of trained agents by optimizing objective functions to encourage policy “diversity” within these small populations. Nevertheless, focusing solely on individual-level behavioral diversity while neglecting the potential impact of population scale may limit the overall effectiveness of population-based methods. In fact, the constrained population size could be one of the critical bottlenecks restricting current ZSC performance.
@@
To address this challenge, we propose a scalable population-based training framework: ScaPT. Figure 1 illustrates the differences among self-play, common population-based training, and our proposed framework. Our goal is to enable ZSC training to scale more efficiently to larger populations, mitigating the constraints imposed by computational resources. To achieve this, our framework is designed to meet two key criteria: (1) alleviate the coupling between population size and computational and parameter overhead, thereby mitigating the linear growth of resource consumption with increasing population size; (2) maintain comparable or even superior ZSC performance relative to conventional population-based methods. (2) maintain comparable or even superior ZSC performance relative to conventional population-based methods.
@@
In experiments conducted on two standard ZSC benchmark environments, matrix games and Hanabi, ScaPT achieves significantly better performance than existing algorithms, primarily due to its scalable population design that allows it to leverage larger agent pools under the same memory resources constraints. Furthermore, when comparing different algorithms under the same scalable population framework, ScaPT demonstrates outstanding zero-shot coordination capabilities.
@@
  • ScaPT designs the scalable population using a hierarchical meta-agent architecture to simulate large populations, achieving a high degree of decoupling between population size and computational resources, while employing conditional mutual information to ensure diversity within the scalable population
@@
Self-play (yu2022surprising) is a commonly-used technique in MARL, being able to quickly train a fixed group of cooperative agents while falling short in zero-shot coordination. To address this issue, some researchers make agents reason about the task or partners. Related frameworks include breaking symmetries of the task to keep agents from learning specified strategies (hu2020other; treutlein2021new; muglich2022equivariant), conducting multi-level reasoning to get higher-level consensus (cui2021k; hu2021off) and requiring agents to predict partners’ actions (lucas2022any; yan2024efficient). The core disadvantage of the above solutions is that agents might form an algorithm-level consensus or overfit to several partners.
Another kind of mainstream solution is population-based training. This solution improves the generalization ability of a main agent by requiring it to cooperate well with all partner agents in a population. Consequently, improving the divergence of the population becomes a primary objective. Typical frameworks include reducing the collaboration scores of different partner agents within the population to make them behave differently (charakorn2022generating; rahman2023generating), improving trajectory diversity of different agents (lupu2021trajectory) and increasing policy entropy of the population (zhao2023maximum). However, this kind of framework may face heavy computational load or limited usage. For example, most population training frameworks train distinct neural network parameters for agents in a population, which is time-consuming; Several frameworks (lupu2021trajectory; zhao2023maximum) require agents to output differentiable action distribution, while some important RL frameworks such as value-based methods may not meet the requirement. In comparison, our proposed ScaPT utilizes a meta-agent to efficiently achieve population training, and designs a generic mutual information term to guarantee population divergence.
@@
In this section, we introduce ScaPT, a scalable population-training framework that enables efficient and diverse zero-shot coordination via a hierarchical meta-agent and a mutual information-based objective.
@@
A major drawback of population training is its high resource demand due to the need to train multiple agents. Considering that some multi-task learning and meta-reinforcement learning study(ma2018modeling; sagirova2025srmt), where meta-networks are trained to quickly adapt to diverse tasks, it is possible to selectively share certain parameters responsible for task-related processing, while employing separate sub-modules to independently optimize behavior-related parameters.
@@
where $H$ represents observation input (containing current observation and historical trajectory used for decision), $U$ represents the index of sub-decision module that outputs action $A$, and $p(a,u,h)$ is the joint probability density function. Considering that reinforcement learning frameworks commonly estimate integrals using the Monte Carlo method, that is, sampling transitions from replay buffer to calculate, the unbiased estimation of mutual information $\hat{I}(A;U|H)$ can be written as:
@@
$-\log\sum_{i=1}^{K}p(u_{i}|h_{j})\cdot p(a_{j}|u_{i},h_{j})\bigg]$
where $K$ is the total number of sub-decision modules, $N$ is the number of transitions, $a_{j},u_{j},h_{j}$ are action, sub-decision module index and observation input of the $j$-th transition, and $p(a_{j}|u_{j},h_{j})$ is the conditional probability that an agent takes action $a_{j}$ given $u_{j},h_{j}$. For brevity, use $I_{j}$ to denote the $j$-th term in $\hat{I}(A;U|H)$:
$I_{j}:=\log p(a_{j}|u_{j},h_{j})-\log\sum_{i=1}^{K}p(u_{i}|h_{j})p(a_{j}|u_{i},h_{j})$ (3)
@@
Computing $\bar{I}(A;U|H)$ requires obtaining all agents’ actions under the same observation, which is efficiently achieved via the meta-agent with a single forward pass—unlike standard population training that needs $N$ slow LSTM passes. Notably, ScaPT’s two components are complementary: the meta-agent enables efficient mutual information computation and training, while the conditional mutual information term guides the learning of a diverse population.
@@ V MPMP, PPPP $\sum_{i=1}^{N}J(\pi_{m},\pi_{pi})$ $J(\pi_{pi},\pi_{pi})+J(\pi_{m},\pi_{pi})$
Use $J(\pi_{1},\pi_{2})$ to denote the expected cumulative discounted return obtained by the collaboration of $\pi_{1}$ and $\pi_{2}$. Use $\pi^{M}_{i}$ to denote the policy obtained with training framework $M$ and random seed $i$. The earliest metric to evaluate zero-shot cooperation performance is intra-algorithm cross-play (abbreviated as Intra-XP) score (hu2020other):
@@
Then, we compare the MEP algorithm under the common and proposed scalable population architecture in computational cost, training time, and ZSC performance on 2-player Hanabi game, to verify that our framework reduces training resource consumption while maintaining comparable performance. As shown in Fig. 4, the scalable population scales more efficiently with slower growth in resource use and training time, while the common framework grows approximately linearly. Both approaches achieve comparable ZSC performance, demonstrating the scalability and effectiveness of our method for complex tasks requiring large population agents.
@@
We conduct a 2-player Hanabi experiment to validate the superiority of ScaPT, in which the population size is set to 5 for all population-based algorithms. As shown in Table 3, the mean and standard error of evaluation metrics are reported. ScaPT achieves the best result on the 1ZSC-XP metric, indicating its strong coordination ability when paired with 40 previously unseen partner algorithms. On the Intra-XP metric, ScaPT attains the highest score among all methods except OBL. Notably, although OBL scores high on Intra-XP, its 1ZSC-XP score is only 3.8, suggesting that OBL lacks genuine zero-shot coordination capability since it performs well mainly when cooperating with agents that are trained using the same algorithm as itself. Considering both metrics together, all population-based methods demonstrate solid performance. Among them, ScaPT delivers the most outstanding results, further confirming its superiority.
@@ I II III IV V VI
MM Score: The self-play score of the main agent;
@@
PP Score: The self-play score of the partner agents;
@@
As is shown in Fig. 6, when the mutual information term $\bar{I}(A;U|H)$ is ignored (i.e. $\alpha=0$), Diff Prob quickly rises to $1$, meaning that the partner agents in the population act similarly. As a result, the generalization performance of the main agent is reduced (low XP scores during testing) despite that the training process goes smoothly (high MM/MP/PP scores during training). When $\alpha$ is set to a proper value ($\alpha=1$), Diff Prob maintains relatively low, indicating a diverse population. When $\alpha$ is set to a large value ($\alpha=10$), the population diversity does not further increase, and what’s worse, the self-play score of partner agents (PP score) goes low. This indicates that the rationality of the partner behavior may be affected due to large $\alpha$, and the coordination performance of the main agent is also hampered. To sum up, $\bar{I}(A;U|H)$ can help build a diverse population and thereby improve the zero-shot coordination, while assigning it a too large weight might have negative effect.
@@ Conclusion
In this paper, we analyzed limitations of population-based ZSC methods and proposed a population-size-free training framework. Experiments demonstrated its effectiveness and revealed a strong correlation between population size and ZSC performance, further validating the significance of our study.
@@ Appendix E Detailed results of 5-player hanabi game with different population-size(ps)
  • Act Group: MM, MP, PP.: There are three kinds of act groups will be used for interacting with the environment and generating transitions: [Main agent, Main agent], [Main agent, Partner agent] and [Partner agent, Partner agent]. Consequently,

### v2

# Efficient Reinforcement Learning for Zero-Shot Coordination in Evolving Games

@@ Abstract
Zero-shot coordination(ZSC), a key challenge in multi-agent game theory, has become a hot topic in reinforcement learning (RL) research recently, especially in complex evolving games. It focuses on the generalization ability of agents, requiring them to coordinate well with collaborators from a diverse, potentially evolving, pool of partners that are not seen before without any fine-tuning. Population-based training, which approximates such an evolving partner pool, has been proven to provide good zero-shot coordination performance; nevertheless, existing methods are limited by computational resources, mainly focusing on optimizing diversity in small populations while neglecting the potential performance gains from scaling population size. To address this issue, this paper proposes the Scalable Population Training (ScaPT), an efficient RL training framework comprising two key components: a meta-agent that efficiently realizes a population by selectively sharing parameters across agents, and a mutual information regularizer that guarantees population diversity. To empirically validate the effectiveness of ScaPT, this paper evaluates it along with representational frameworks in Hanabi cooperative game and confirms its superiority.
@@ Introduction
As Multi-Agent Reinforcement Learning (MARL) continues to achieve remarkable success in a wide range of complex games and domains, its scalability is increasingly hindered by the greedy growth of computational resource demands, in the scenarios like agent MOBA games (gao2023towards), robotic soccer (zhu2024dynamic), and intelligent manufacturing systems (do2025heterogeneous) owing to the demand for training a diverse and evolving set of agents. These challenges highlight the need for more scalable and resource-efficient MARL frameworks.
@@
We address this problem in the context of Zero-Shot Coordination (ZSC)—a setting where agents must collaborate with unseen partners. This scenario is central to what we term “evolving games”: environments where the pool of potential partners is not fixed, but rather represents a continuously shifting or evolving set of strategies. While self-play (lowe2017multi) is widely used for training cooperation, it often leads to overfitting, limiting agents’ ability to generalize. To overcome this, hu2020other introduced the ZSC problem, where agents are evaluated on their ability to coordinate without prior exposure to their partners. Population-based training (jaderberg2017population) offers a promising solution by exposing agents to a diverse set of behaviors during training, thus encouraging more generalizable coordination strategies (charakorn2023generating). In principle, if the training population covers all potential partner strategies, ZSC would be effectively solved. However, approaching this ideal requires training a large and diverse population of agents, which is itself constrained by limited computational resources, hindering the development of efficient and scalable RL methods.
Figure 1: The diagram of different training paradigms in evolving games.
@@
Recent methods focus on improving Zero-Shot Coordination (ZSC) performance by enhancing population diversity, which serves as a key proxy for strategic variation in these evolving games. For example, Trajectory Diversity for Zero-Shot Coordination (lupu2021trajectory) maximizes trajectory-level diversity to foster generalizable strategies. Similarly, Maximum Entropy Population-Based Training (zhao2023maximum) uses entropy maximization to broaden behavioral coverage and improve human-AI coordination. Heterogeneous Multi-Agent Zero-Shot Coordination by Coevolution (xue2024heterogeneous) tackles real-world heterogeneity by co-evolving two distinct populations through iterative pairing and selection, enabling robust cooperation with diverse partners.
@@
However, constrained by limited computational resources, existing studies typically maintain relatively small population sizes and attempt to improve the zero-shot coordination performance of trained agents by optimizing objective functions to encourage policy “diversity” within these small populations. Nevertheless, focusing solely on individual-level behavioral diversity while neglecting the potential impact of population scale may limit the overall effectiveness of population-based methods in truly capturing the breadth of an evolving game’s strategy space. In fact, the constrained population size could be one of the critical bottlenecks restricting current ZSC performance.
@@
To address this challenge, we propose an efficient RL with scalable population-based training framework: ScaPT. Figure 1 illustrates the differences among self-play, common population-based training, and our proposed framework in evolving games. Our goal is to enable ZSC training to scale more efficiently to larger populations, mitigating the constraints imposed by computational resources. To achieve this, our framework is designed to meet two key criteria: (1) alleviate the coupling between population size and computational and parameter overhead, thereby mitigating the linear growth of resource consumption with increasing population size; (2) maintain comparable or even superior ZSC performance relative to conventional population-based methods. (2) maintain comparable or even superior ZSC performance relative to conventional population-based methods.
@@
In experiments conducted on two standard ZSC benchmark environments, matrix games and Hanabi, ScaPT achieves significantly better performance than existing algorithms, primarily due to its scalable population design that allows it to leverage larger agent pools under the same memory resources constraints. Furthermore, when comparing different algorithms under the same scalable population framework, ScaPT demonstrates outstanding zero-shot coordination game capabilities.
@@
  • ScaPT designs the scalable population using a hierarchical meta-agent architecture to simulate large populations, achieving a high degree of decoupling between population size and computational resources, while employing conditional mutual information to ensure diversity within the scalable population.
@@
Self-play (yu2022surprising) is widely used in MARL, and can efficiently train cooperative agents, but it typically fails under zero-shot coordination. To address this issue, recent works introduce additional reasoning abilities, such as break symmetries of the task to keep agents from learning specified strategies (hu2020other; treutlein2021new; muglich2022equivariant), conduct multi-level reasoning for higher-level consensus (cui2021k; hu2021off) and predict partners’ actions (lucas2022any; yan2024efficient). However, these approaches often lead to algorithm-level consensus or overfitting to a limited set of partners.
Population-based training enhances generalization by exposing a main agent to diverse partners. Common approaches include reducing the collaboration scores of different partner agents within the population to make them behave differently (charakorn2023generating; rahman2023generating), improving trajectory diversity of agents (lupu2021trajectory) and increasing policy entropy of the population (zhao2023maximum). However, most approaches train distinct networks per agent or require differentiable action distributions (lupu2021trajectory; zhao2023maximum), limiting efficiency and applicability; In contrast, ScaPT employs a meta-agent for efficient population training, and designs a generic mutual information term to guarantee population divergence. This differs from MI-based skill-discovery methods such as DIAYN (eysenbach2018diversity) and DADS (sharma2019dynamics), which increase intra-agent skill diversity rather than inter-agent policy diversity.
@@
We present ScaPT, a scalable population-training framework that achieves diverse zero-shot coordination.
@@
A major drawback of population training is its high resource demand. Inspired by meta-RL and multi-task learning (ma2018modeling; sagirova2025srmt), task-related parameters can be shared, while behavior-specific submodules are optimized independently.
@@
where $H$ represents observation input (containing current observation and historical trajectory used for decision), $U$ represents the index of sub-decision module that outputs action AA, and $p(a,u,h)$ is the joint probability density function. Considering that reinforcement learning frameworks commonly estimate integrals using the Monte Carlo method, that is, sampling transitions from replay buffer to calculate, the unbiased estimation of mutual information $\hat{I}(A;U|H)$ can be written as:
@@
$\qquad-\log\sum_{i=1}^{K}p(u_{i}|h_{j})\cdot p(a_{j}|u_{i},h_{j})\bigg]$
where $K$ is the total number of sub-decision modules, NN is the number of transitions, $a_{j},u_{j},h_{j}$ are action, sub-decision module index and observation input of the $j$-th transition, and $p(a_{j}|u_{j},h_{j})$ is the conditional probability that an agent takes action $a_{j}$ given $u_{j},h_{j}$. For brevity, use $I_{j}$ to denote the $j$-th term in $\hat{I}(A;U|H)$. The equation is: $I_{j}:=\log p(a_{j}|u_{j},h_{j})-\log\sum_{i=1}^{K}p(u_{i}|h_{j})p(a_{j}|u_{i},h_{j})$.
@@
Computing $\bar{I}(A;U|H)$ requires obtaining all agents’ actions under the same observation, which is efficiently achieved via the meta-agent with a single forward pass—unlike standard population training that needs NN slow LSTM passes. Notably, ScaPT’s two components are complementary: the meta-agent enables efficient mutual information computation and training, while the conditional mutual information term guides the learning of a diverse population.
@@ V MPMP, PPPP $\sum_{i=1}^{N}J(\pi_{m},\pi_{pi})$ $J(\pi_{pi},\pi_{pi})+J(\pi_{m},\pi_{pi})$
Use $J(\pi_{1},\pi_{2})$ to denote the expected cumulative discounted return obtained by the collaboration of $\pi_{1}$ and $\pi_{2}$. Use $\pi^{M}_{i}$ to denote the policy obtained with training framework $M$ and random seed ii. The earliest metric to evaluate zero-shot cooperation performance is intra-algorithm cross-play (abbreviated as Intra-XP) score (hu2020other):
@@
To validate our archictecture and ensure fair comparison, we evaluate MEP under common and scalable population architectures in computational cost, training time, and ZSC performance on 2-player Hanabi. As shown in Fig. 4, the scalable population scales more efficiently with slower growth in resource use and training time, while the common framework grows approximately linearly. Both approaches achieve comparable ZSC performance, demonstrating the scalability and effectiveness of our method for complex tasks requiring large population agents.
@@
We conduct a 2-player Hanabi experiment with population size 5 for all population-based algorithms. Table 3 reports mean and standard error of evaluation metrics. ScaPT achieves the highest 1ZSC-XP, demonstrating strong coordination with 40 unseen partners and ranks top on Intra-XP except for OBL. Although OBL scores high on Intra-XP, its 1ZSC-XP score is only 3.8, suggesting that OBL lacks genuine zero-shot coordination capability since it performs well mainly when cooperating with agents that are trained using the same algorithm as itself. Considering both metrics together, all population-based methods demonstrate solid performance. Among them, ScaPT delivers the most outstanding results, further confirming its superiority.
@@ I II III IV V VI
$M$ Score: The self-play score of the main agent;
@@
$P$ Score: The self-play score of the partner agents;
@@
As is shown in Fig. 6, when the mutual information term $\bar{I}(A;U|H)$ is ignored (i.e. $\alpha=0$), Diff Prob quickly rises to 11, meaning that the partner agents in the population act similarly. As a result, the generalization performance of the main agent is reduced (low XP scores during testing) despite that the training process goes smoothly (high MM/MP/$P$ scores during training). When $\alpha$ is set to a proper value ($\alpha=1$), Diff Prob maintains relatively low, indicating a diverse population. A larger $\alpha$ ($\alpha=10$) does not further increase diversity and reduces partner self-play performance ($P$ score), suggesting that excessive weight can harm rationality and coordination. To sum up, $\bar{I}(A;U|H)$ can help build a diverse population and thereby improve the zero-shot coordination, while assigning it a too large weight might have negative effect.
@@ Conclusion
In this paper, we analyzed limitations of population-based RL methods for ZSC, which struggle to scale efficiently enough to represent the diversity of evolving games, and proposed a population-size-free RL training framework. Experiments demonstrated its effectiveness and revealed a strong correlation between population size and ZSC performance, further validating the significance of our study.
@@ Appendix E Detailed results of 5-player hanabi game with different population-size(ps)
  • Act Group: $M$, MP, $P$.: There are three kinds of act groups will be used for interacting with the environment and generating transitions: [Main agent, Main agent], [Main agent, Partner agent] and [Partner agent, Partner agent]. Consequently,
@@
Acknowledgments
This work is supported by National Key Research and Development Program of China (under Grant No.2023YFB2903904), and BNRist projects (No.BNR20231880004 and No.BNR2024TD03003).